# Motion-compensated cardiac MRI using low-rank diffeomorphic flow (DMoCo)

### v2

@@ V-C Human subjects
The in-vivo data was acquired from five cardiac patients at the University of Iowa Hospitals and Clinics (UIHC), according to the institutional review board (IRB) protocols and the consent of the subject. The proposed sequence was added to the end of an existing clinical protocol, approximately 25 to 35 minutes after the short-acting Gadolinium (Gd) contrast agent was administered intravenously.
For qualitative comparisons, we extracted the time series of slices from the 3D reconstructions that closely match the 2D breath-held acquisitions with different views (e.g., short-axis, four-chamber, and two-chamber views). We note that separate breath-holds are required for the acquisition of each of the 2D stacks, while the proposed approach only involves a single free-breathing 3D acquisition. We note that the image contrast of the 3D acquisitions is drastically different from that of the 2D acquisitions, making quantitative image comparisons challenging.
Figure 3: Impact of the velocity constraint: We show the recovered images and the deformation maps using (a) the proposed model, (b) the proposed model without the velocity constraint, and the MoCo-SToRM method, which directly models the deformation maps rather than integrating the velocities. The $l_{2}$ norm of the deformations is reported under each column denoted (MD); a higher number indicates more complex deformation. The results show that the velocity constraint necessarily results in smoother deformations than without it. We also note that the proposed DMoCo model is more constrained than the MoCo-SToRM approach.
@@ VI-A Impact of low-rank approximation
Figure 4: We used the XCAT framework to simulate a free-breathing and ungated acquisition with the same settings as the in-vivo acquisition. We corrupted the simulated k-space observations with Gaussian noise of standard deviation $5\times 10^{-3}$. We note that the acquisition scheme corresponds to a highly undersampling of several motion phases (e.g., systolic phases in end inspiration). The top three rows show a failure case, where the MoCo-SToRM approach resulted in a discontinuity in the recovered time profiles, shown on the right. This can also be appreciated from the recovered images (left panel with a blue border, corresponding to the blue lines in the time profiles on the right panel) in the corresponding phase, which shows increased blurring and a lower PSNR. By contrast, the DMoCo approach with a more constrained deformation model offers an improved reconstruction of these phases (see yellow arrows). The bottom rows correspond to reconstructions from the systolic and diastolic phases, indicated by the light blue and green lines in the time profile. We observe increased noise blurring in the MoCo-SToRM reconstructions compared to DMoCo.
We first study the low-rank nature of the velocity and deformation tensors in Figure 2 in a 2D annulus phantom in left column (a).
@@
The upper rows show three images at different physiological phases, with the template/source in the systolic phase and the target. The middle image corresponds to an intermediate phase. We perform a joint registration of volumes in the image domain using (13). We overlay the deformation plots on the template in the middle row, while the velocity plots are overlaid in the bottom row. We observe that the velocity fields are roughly similar across all phases, while the deformations are significantly different; the phase-dependent deformations can be obtained by integrating a relatively phase-independent velocity field according to (13). We note that the integration in (6) relies on the velocity fields in the deformed state and therefore are non-linearly dependent on the velocity fields. In the bottom row 2, we measure the distortion of the deformed images in each phase from the respective reference image as a function of the rank of the velocity (orange), deformation fields (blue), and image series (green). Here, we use a 2D annulus with both respiratory and cardiac motion. We estimated the velocity and the corresponding deformations from the reference images using (13). We then (a) truncated the velocity fields to different ranks, and evaluated the corresponding deformation fields by integrating the approximated velocity fields as in (6), and (b) performed an SVD of the deformation fields at all phases and truncated to a lower rank model. We compare the warped images with the true ones to generate the percent error plots.
@@
The figures show that a rank=3 model on the velocity is sufficient to yield low MSE values 3.45%; converts to 22.3%, while even a rank=10 approximation of the deformation fields can not give the same error rate as a rank=3 model using diffeomorphic flow, 3.45% to 10.2%. The compact velocity representation is expected to yield more robust motion estimates, especially when estimated from noisy and undersampled images. This is considered in the subsequent experiments.
@@ VI-B Impact of velocity constraint
We study the impact of the velocity constraint in (11) in Fig. 3. Here, we used (13) for the joint fit in the image domain, with the end systolic frame as the template. We show the estimated deformations for different cardiac phases, estimated using MoCo-SToRM in (a), the proposed model with the velocity constraint in (b), and without the velocity constraint in (c). We note that all models result in deformed images that closely match the original ones, evidenced by the metrics Fig. 3 table 1. This is expected because noise-free fully sampled images are used for image domain fitting. The results show that without the velocity constraint, the proposed scheme produces highly oscillatory deformations, indicated by the high $\ell_{2}$ norm of the deformation in (c); we note that there are numerous deformations of the template that can yield the targets. The addition of the velocity constraint regularizes the deformations, resulting in deformations with significantly lower $\ell_{2}$ norm in (b). We also note that the deformations offered by MoCo-SToRM in (a), which directly models the deformation fields rather than the velocity fields, result in deformations that are more complex than the proposed scheme. These results show that DMoCo with velocity constraints offer a constrained model that can account for the motion, while yielding smoother deformations. The resulting algorithm is thus expected to be less sensitive to undersampling artifacts and noise, which is studied in the next subsection.
@@
We compare the MoCo-SToRM and DMoCo results on the X-CAT phantom in Fig. 4, where the reconstruction is performed from noisy measurements in k-space. We considered the same sampling pattern as in the in-vivo experiments. We note from the time-profile plots that the MoCo-SToRM results in a piecewise-constant profile with abrupt jumps. We attribute the rapid jumps in the MoCo-SToRM temporal profiles to the less constrained deformation model. By contrast, the DMoCo approach yields smoother profiles, which are consistent with the gold standard. The main distinction between the two approaches is the way in which the deformations are generated. The low rank nature of the velocities that are integrated to obtain the deformations as in (6) and the velocity constraints within DMoCo encourages the deformations to be a smooth function of the cardiac phases. We also observe that the images recovered using DMoCo are less blurred and more consistent with the reference compared to MoCo-SToRM. Finally, we measure the respiratory displacement from the left dome of the diaphragm to be $26.06$ mm for DMoCo, $21.69$ mm for MoCo-SToRM, compared to the ground truth of $27.33$ mm for the XCAT phantom.
@@ VI-D In Vivo Results
The reconstructed images using the proposed DMoCo approach, MoCo-SToRM, and a motion resolved approach from two patients are shown in Fig. 5. The first row corresponds to short-axis reconstructions in the systole, while the second row corresponds to the diastole. We also show results in the two-chamber and four-chamber views. The first column denotes the 2D SSFP cine acquisitions. We note that the contrast of the 2D cine acquisition is superior to that of the 3D acquisition, mainly because of the inflow effects. In contrast, the saturation of the magnetization in the 3D radial acquisition translates to a lower contrast-to-noise ratio between the blood pool and the myocardium. This is consistent with the literature [20].
We note that the reconstructions with motion-resolved (MoR) method appear more noisy than the motion-compensated (MoCo) schemes. As noted before, the motion-resolved schemes are sensitive to uneven sampling of the bins; the radial undersampling artifacts manifest as noise in the reconstructions. In contrast, motion-compensated methods combine the data from different motion states to recover the static template, which translates to reducing noise and artifacts. We do not show the temporal profiles of MR because it only has five cardiac phases. The results also show that the time profiles of the proposed model are more comparable to 2D cine, compared to MoCo-SToRM, which tends to result in more averaged time profiles.
Additionally, 3D bSSFP lacks contrast due to blood flowing into the imaging plane like 2D cine. Interestingly, we found the static template converged to ventricular systole even though the heart spends less time in systole than diastole.
@@ VII Discussion
The proposed DMoCo approach represents the images at the different motion phases as deformed versions of a static template. The deformation at a specific motion phase is represented as a diffeomorphism, obtained by integrating a parametric velocity field over intermediate phases as shown in (7). The parametric velocity field itself is represented using a low-rank model, while a velocity constraint is added to encourage the deformations to be path-independent. The parameters of the low-rank model and the static template are learned directly from the k-space data in a patient-specific, self-supervised fashion. The experiments on the numerical phantom show that the DMoCo model is more constrained than the MoCo-SToRM approach, which uses a low-rank representation to directly model the deformations. In particular, Fig. 2 shows that the velocity fields, which are integrated to produce the deformation fields, can be of a lower rank than the deformation fields themselves. We also see from the results in Fig. 3 that the addition of the velocity constraints further regularizes the representation, while being able to accurately model the deformations. The integration of velocity fields over the different intermediate phases to produce the deformation serves as a smoothness constraint between adjacent motion phases, translating to improved preservation of time profiles, as seen from Fig. 4; the lack of a smoothness constraint in MoCo-SToRM resulted in abrupt discontinuities in the time profiles and blurred reconstructions. This trend is also supported by the respiratory displacement (RD) which is $21.69$mm for MoCo-SToRM and $26.06$mm for DMoCo compared to the ground truth of $27.33$mm.
The experiments on in-vivo data are in line with the phantom results. We observe that the time profiles of DMoCo are more in agreement with the 2D cine acquisitions. The proposed scheme reduced the noise and alias artifacts compared to a motion-resolved reconstruction; the averaging of the different motion phases to yield the template enables the motion compensation methods to be more resilient to undersampling and noise.
@@
Despite promising results, the proposed study still has some limitations that restrict data quality, which we aim to overcome in our future studies. A challenge is the low contrast-to-noise ratio (CNR) in 3D acquisitions. Unlike 2D scans that present good inflow contrast from unsaturated blood that flows into the imaging slice, the contrast in 3D acquisitions is inherently low. A possible alternative is a 3D multislab or interleaved multi-slice acquisition, which may offer improved contrast. Another challenge is that our current sequence did not employ fat suppression, resulting in strong residual streak artifacts. We will employ intermittent pulses of fat saturation as in [5] to minimize residual fat signals. Furthermore, current studies show the exciting possibility of significantly improving CNR in 3D studies by using ferumoxytol [4], which may be another option to improve CNR.
The computational complexity of the current unsupervised approach is around 3 hours to reconstruct a dataset on an A100 GPU. While this is faster than the current motion resolved approaches, we believe that it can be accelerated in the future. We will consider more efficient variable splitting strategies or improved deep learning based priors (e.g. [7]) on the template or initialization strategies, which may help manage the computational complexity.
@@ VIII Conclusion
An unsupervised motion-compensated image reconstruction algorithm was introduced for free-breathing and ungated 3D cardiac MRI. The images at different motion phases were expressed as the deformation of a static image template, where the deformations are represented using low-rank diffeomorphic flow model. The static template and low-rank motion model parameters were learned directly from the k-space data in an unsupervised manner. The experiments show that the MoCo approach is less sensitive to undersampling of the motion phases. The more constrained motion model translated to more robust motion estimates, which offers improved performance compared to MoCo methods that directly model the deformations.

### v3

@@ V-C Human subjects
The in-vivo data was acquired from five cardiac patients at the University of Iowa Hospitals and Clinics (UIHC), according to the institutional review board (IRB) protocols and the consent of the subject. The proposed sequence was added to the end of an existing clinical protocol, approximately 15 to 20 minutes after the short-acting Gadolinium (Gd) contrast agent was administered intravenously $0.1\leavevmode\nobreak\ mmol/kg$.
For qualitative comparisons, we extracted the time series of slices from the 3D reconstructions that closely match the 2D breath-held acquisitions with different views (e.g., short-axis, four-chamber, and two-chamber views). We note that separate breath-holds are required for the acquisition of each of the 2D stacks, while the proposed approach only involves a single free-breathing 3D acquisition. We note that the image contrast of the 3D acquisitions is drastically different from that of the 2D acquisitions, making quantitative image comparisons challenging. We compare the ejection fraction (EF), end-systolic volume (ESV), and end-diastolic volume (EDV) against the values found in the 2D cine.
Figure 3: Impact of velocity constraint: We show the registered images and the deformation maps using (a) the proposed model, (b) the proposed model without the velocity constraint, and (c) the MoCo-SToRM method, which directly represents the deformation maps rather than integrating the velocities. The $l_{2}$ norm of the deformations is reported under each column denoted (MD); a higher number indicates more complex deformation. We also show the PSNR and SSIM measures of the moved images using all the methods, which show that all of them offer good registration. We note that registration is an ill-posed problem with multiple feasible deformations resulting in the same image. Our results show that the use of the velocity constraint translates to smoother deformations (lower MD and less wiggly deformations) than without the velocity constraint. We also note that the proposed DMoCo model is also more constrained than the MoCo-SToRM approach.
@@ VI-A Impact of low-rank approximation
Figure 4: Comparison using XCAT phantom: We used the XCAT framework to simulate a free-breathing and ungated acquisition with the same settings as the in-vivo acquisition. We corrupted the simulated k-space observations with Gaussian noise of standard deviation $5\times 10^{-3}$. We note that the acquisition scheme corresponds to a highly undersampling of several motion phases (e.g., systolic phases in end inspiration). The top three rows show a failure case, where the MoCo-SToRM approach resulted in a discontinuity in the recovered time profiles, shown on the right. This can also be appreciated from the recovered images (left panel with a blue border, corresponding to the blue lines in the time profiles on the right panel) in the corresponding phase, which shows increased blurring and a lower PSNR. By contrast, the DMoCo approach with a more constrained deformation model offers an improved reconstruction of these phases (see yellow arrows). The bottom rows correspond to reconstructions from the systolic and diastolic phases, indicated by the light blue and green lines in the time profile. We observe increased noise blurring in the MoCo-SToRM reconstructions compared to DMoCo.
@@
We first study the low-rank nature of the velocity and deformation tensors in Figure 2 in a 2D annulus phantom in the left column (a). The upper rows show three images at different physiological phases, with the right-most image as the template/source, which is in the systolic phase. The middle image corresponds to an intermediate phase. We perform a joint registration of volumes in the image domain using (13). We overlay the deformation plots on the template in the middle row, while the velocity plots are overlaid in the bottom row. We observe that the velocity tensors are roughly similar across all phases, while the deformations are significantly different; the phase-dependent deformations can be obtained by integrating a relatively phase-independent velocity tensor according to (13). We note that the integration in (6) relies on the velocity tensors in the deformed state and therefore are non-linearly dependent on the velocity tensors. In the bottom row 2, we measure the distortion of the deformed images in each phase from the respective reference image as a function of the rank of velocity (orange), deformation fields (blue), and image series (green). Here, we use a 2D annulus with both respiratory and cardiac motion. We estimated the velocity and the corresponding deformations from the reference images using (13). We then (a) truncated the velocity tensors to different ranks, and evaluated the corresponding deformation fields by integrating the approximated velocity tensors as in (6), followed by deforming the template to the respective phases using the corresponding deformation fields, (b) performed a low rank approximation of the deformation fields, followed by deforming the template to the respective phases using the low-rank approximations, and (c) performed a low-rank approximation of the images in the time series. We compare the approximate images to the actual ones to generate the error plot.
@@
The figures show that a velocity model of rank = 3 in the velocity is sufficient to produce low MSE values of 3. 45%, while even a rank=10 approximation of the deformation fields and images can only give errors of 10.2% and 22.3%, respectively. The compact velocity representation is expected to yield more robust motion estimates, especially when estimated from noisy and undersampled images. This is considered in subsequent experiments.
@@ VI-B Impact of velocity constraint
We study the impact of the velocity constraint in (11) in Fig. 3. We set the rank of the velocity model to ten. Here, we used (13) for the joint fit in the image domain, with the end systolic frame as the template. We show the estimated deformations for different cardiac phases, estimated using the proposed model with the velocity constraint in (a), without the velocity constraint in (b), and MoCo-SToRM in (c). We note that all models result in deformed images that closely match the original ones, as evidenced by the PSNR and SSIM metrics reported in the table in Fig. 3. We note that there are numerous deformations of the template that can yield the targets. The results show that without the velocity constraint, the proposed scheme in (b) results in oscillatory deformations, indicated by the higher $\ell_{2}$ norm of the deformation (MD) shown below the images. The addition of the velocity constraint regularizes the deformations, resulting in deformations with significantly lower MD norm in (a). We also note that the deformations offered by MoCo-SToRM in (c), which directly models the deformation fields rather than the velocity tensors, result in deformations that are more complex than those of the proposed scheme. We note that the deformations are significantly more oscillatory in regions with low intensity away from the heart that are not seen from the above figures; these are not shown due to space constraints. These results show that DMoCo with velocity constraints offers a more constrained model that can account for motion while yielding smoother deformations. The resulting algorithm is thus expected to be less sensitive to undersampling artifacts and noise, which is studied in the next subsection.
@@
We compare MoCo-SToRM and DMoCo on the X-CAT phantom in Fig. 4, where the reconstruction is performed from noisy measurements in k-space. We considered the same sampling pattern as in the in-vivo experiments. We note from the time-profile plots that the MoCo-SToRM results in a piecewise-smooth profile with abrupt jumps, compared to the reference profiles. We attribute the rapid jumps in the MoCo-SToRM temporal profiles to the less constrained deformation model. In contrast, the DMoCo approach yields smoother profiles, which are more consistent with the gold standard. The main distinction between the two approaches is the way in which the deformations are generated. The low-rank nature of the velocity tensors, which are integrated to obtain the deformations according to (6), as well as the velocity constraints within DMoCo encourage the deformations to be a smooth function of the cardiac phases. We also observe that the images recovered using DMoCo are less blurred and more consistent with the reference compared to MoCo-SToRM. Finally, we measure the respiratory displacement from the left dome of the diaphragm to be $26.06$ mm for DMoCo, $21.69$ mm for MoCo-SToRM, compared to the ground truth of $27.33$ mm for the XCAT phantom, which also illustrates the improved accuracy of respiratory motion estimates.
@@ VI-D In Vivo Results
The reconstructed images using the proposed DMoCo approach, MoCo-SToRM, and a motion resolved approach from two patients are shown in Fig. 5. The first row of each patient corresponds to short-axis reconstructions in the systole, while the second row corresponds to the diastole. We also show results in the two-chamber and four-chamber views. The first column denotes the 2D SSFP cine acquisitions. We note that the contrast of the 2D cine acquisition is superior to that of the 3D acquisition. This is due to the inflow of unsaturated blood out of the slice to the slice, which appears significantly brighter than the saturated myocardium that experiences repeated radiofrequency excitation pulses. In contrast, the entire blood pool along with the myocardium is repeatedly excited in 3D acquisitions, translating to somewhat similar saturation levels. This results in a lower contrast-to-noise ratio between the blood pool and the myocardium. This is consistent with the literature [33].
We note that the reconstructions with motion-resolved (MoR) approach appear more noisy than the motion-compensated (MoCo) schemes. As noted before, the motion-resolved schemes are sensitive to uneven sampling of the bins; the radial undersampling artifacts manifest as noise in the reconstructions. In contrast, motion-compensated methods combine the data from different motion states to recover the static template, which translates to reducing noise and artifacts. We do not show the temporal profiles of the MoR because it has only five cardiac phases. The results also show that the time profiles of the proposed model are more comparable to 2D cine, compared to MoCo-SToRM, which tends to have flatter time profiles.
We segmented the left ventricles of the DMoCo reconstructions as shown in Fig. 6, which we compare against the 2D cine segmentations also shown. We compare the end diastolic and end systolic volumes, as well as the ejection fraction, in Fig. 7. The results show moderate agreement between the 3D and 2D measures. We note that DMoCo offers isotropic resolution, whereas 2D cine offers highly anisotropic volumes and thick slices. These differences often translates to inconsistencies between the measures, which is also observed in several previous studies that compare functional measures between 3D and 2D [34, 35, 5].
Figure 6: Reconstructed volumes and segmentations from patient three: We exported the short-axis projections of the 3D data from DMoCo reconstructions. The endocardium in each of the projections were segmented from both 2D cine and 3D DMoCo reconstructions using Segment-It, a free software from MEDVISE [36]. (b) indicates the DMoCo reconstructions in end systole and diastole, together with the LV segmentations. (c) are the corresponding 2D BH cine slices, together with their segmentations. (a) is a 3D visualization if the segmented volumes using each method.
Figure 7: Comparison of functional measures with 2D BH cine acquisitions: We compared end-systolic, end-diastolic, and ejection fractions from DMoCo to 2D cine. We also show best fit line that passes through the origin.
@@ VII Discussion
The proposed DMoCo approach represents the images in the different motion phases as deformed versions of a static template. The deformation in each specific motion phase is represented as a diffeomorphism, obtained by integrating a parametric velocity tensor over intermediate phases, as shown in (7). The parametric velocity tensor itself is represented using a low-rank model, while a velocity constraint is added to encourage the deformations to be path independent. The parameters of the low-rank model and the static template are learned directly from the k-space data in a patient-specific, self-supervised fashion.
The experiments on the numerical phantom show that the DMoCo model is more constrained than the MoCo-SToRM approach that uses a low-rank representation to directly model the deformations. Specifically, as illustrated in Fig. 2, the velocity model with a significantly lower rank can achieve high-quality deformations, compared to directly modeling the deformations or images using a low-rank representation. We also see from the results in Fig. 3 that the addition of the velocity constraints further regularizes the representation, while being able to accurately model the deformations. The integration of velocity tensors over the different intermediate phases to produce the deformation serves as a smoothness constraint between adjacent motion phases, compared to MoCo-SToRM. We see from Fig. 4 that the lack of smoothness constraints in MoCo-SToRM resulted in abrupt discontinuities in the time profiles and blurred reconstructions. In contrast, the more constrained DMoCo deformation model yielded better preservation of the time profiles. This trend is also supported by more accurate measures of respiratory displacement (RD).
The experiments on in-vivo data are in line with the phantom results. We observe that the time profiles of DMoCo are more in agreement with the 2D cine acquisitions. The proposed scheme reduced the noise and alias artifacts compared to a motion-resolved reconstruction. The combination of information from different motion phases to reconsruct the template enables MoCo methods to be more resilient to undersampling and noise. The ventricular volumes and ejection fractions measured from the 3D data are in moderate agreement with those obtained from the 2D cine measurements, which is consistent with previous studies [34, 35, 5]. A study in a larger cohort of patients is needed for a more thorough comparison.
@@
Despite promising results, the proposed study still has some limitations that restrict data quality, which we aim to overcome in our future studies. A challenge is the low contrast-to-noise ratio (CNR) of the 3D acquisitions. Unlike 2D scans that present good inflow contrast from unsaturated blood that flows into the imaging slice, the contrast in 3D acquisitions is inherently low. A possible alternative is a 3D multislab or interleaved multi-slice acquisition, which may offer improved contrast. Another challenge is that our current sequence did not employ fat suppression, resulting in strong residual streak artifacts. We will employ intermittent pulses of fat saturation as in [37] to minimize residual fat signals. Furthermore, current studies show the exciting possibility of significantly improving CNR in 3D studies by using Ferumoxytol [38, 39, 10], which may be another option to improve CNR.
The computational complexity of the current unsupervised approach is around 3 hours to reconstruct a dataset on an A100 GPU. While this is faster than the current motion resolved approaches, we believe that it can be accelerated in the future. We will consider more efficient variable splitting strategies or improved deep learning based priors (e.g. [40]) on the template or initialization strategies, which may help manage the computational complexity. This will be the focus of our further research. The main focus of the current work is to introduce a novel method, and demonstrate its feasibility in a small group of subjects. A more elaborate study with more subjects is needed to fully evaluate the potential of this method in the clinic.
@@ VIII Conclusion
An unsupervised motion-compensated image reconstruction algorithm was introduced for free-breathing and ungated 3D cardiac MRI. The images in different motion phases were expressed as the deformation of a static image template, where the deformations are represented by a low-rank diffeomorphic flow model. The static template and low-rank motion model parameters were learned directly from the k-space data in an unsupervised manner. The experiments show that the MoCo approach is less sensitive to undersampling of the motion phases. The more constrained motion model translated to more robust motion estimates, which offers improved performance compared to MoCo methods that directly model the deformations.